# Stability of a passive viscous droplet in a confined active nematic liquid crystal

### v1

@@ V Conclusions
We have analyzed the linear stability of a passive viscous Newtonian drop immersed in an active nematic liquid crystal in two dimensions. In extensile systems, activity was found to stabilize the translational mode, but to destabilize deformation modes for small to intermediate drop sizes; these findings are reversed in contractile systems. Elasticity has a predominantly stabilizing effect, except in the case of small drops with low viscosity ratios. We also considered the inverted problem of an active nematic drop surrounded by a viscous layer, where extensile stresses were found to be always destabilizing. When active, elastic, and capillary stresses act in concert, complex trends are predicted, with a non-trivial and non-intuitive dependence of the stability results on droplet size and viscosity ratio.
Our model relied on several strong assumptions that may need to be relaxed or improved upon for comparison with specific experimental systems. Specifically, we considered a sharply aligned liquid crystal in the limit of strong nematic relaxation ($\mathrm{Pe}\rightarrow 0$) and with strong anchoring boundary conditions. A consequence of these choices is that the configuration of the nematic is fully determined by the instantaneous geometry of the system and is not directly affected by the viscous flow driven by the various stresses in the bulk of the nematic. If the assumption of strong elasticity is relaxed, additional instabilities can arise that involve spontaneous swirling flows, as has been observed in experiments Woodhouse and Golstein (2012); Ray et al. (2023) and other models Fürthauer et al. (2012); Theillard et al. (2017); the impact of such flows on the stability problem analyzed here in unknown. The potential emergence and role of topological defects is another question of interest: no defects were present in the model analyzed here (except for the $+1$ defect at the origin in the inverted problem of Sec. IV), but defects may arise in other geometries Chandler and Spagnolie (2024); Neville et al. (2024) and are likely to have an impact on the stability and dynamics. The interaction of multiple droplets suspended in an active nematic is also an open topic of interest.
@@ Acknowledgements.
The authors acknowledge financial support from National Science Foundation Grant No. DMS-2153520 (D.S.).
@@
$\frac{\text{d}^{4}G_{m}}{\text{d}r^{4}}+\frac{2}{r}\frac{\text{d}^{3}G_{m}}{\text{d}r^{3}}-\frac{2m^{2}+1}{r^{2}}\frac{\text{d}^{2}G_{m}}{\text{d}r^{2}}+\frac{2m^{2}+1}{r^{3}}\frac{\text{d}G_{m}}{\text{d}r}+\frac{m^{4}-4m^{2}}{r^{2}}=0,$ (A-1)

### v2

@@ V Conclusions
We have analyzed the linear stability of a passive viscous Newtonian drop immersed in an active nematic liquid crystal in two dimensions. In extensile systems, activity was found to stabilize the translational mode, but to destabilize deformation modes for small to intermediate drop sizes; these findings are reversed in contractile systems. Elasticity has a predominantly stabilizing effect, except in the translational mode where it destabilizes drops of small to intermediate sizes. We also considered the inverted problem of an active nematic drop surrounded by a viscous layer, where extensile stresses were found to be always destabilizing. When active, elastic, and capillary stresses act in concert, complex trends are predicted, with a non-trivial and non-intuitive dependence of the stability results on droplet size and viscosity ratio. Note that we have analyzed the translation and deformation modes separately: in an experimental system, multiple modes will typically be excited simultaneously, and the mode subject to the largest unstable growth rate is expected to dominate the dynamics.
Our model relied on several strong assumptions that may need to be relaxed or improved upon for comparison with specific experimental systems. Specifically, we considered a sharply aligned liquid crystal in the limit of strong nematic relaxation ($\mathrm{Pe}\rightarrow 0$) and with strong anchoring boundary conditions. A consequence of these choices is that the configuration of the nematic is fully determined by the instantaneous geometry of the system and is not directly affected by the viscous flow driven by the various stresses in the bulk of the nematic. If the assumption of strong elasticity is relaxed, additional instabilities can arise that involve spontaneous swirling flows, as has been observed in experiments Woodhouse and Goldstein (2012); Ray et al. (2023) and other models Fürthauer et al. (2012); Theillard et al. (2017); the impact of such flows on the stability problem analyzed here in unclear. The potential emergence and role of topological defects is another question of interest: no defects were present in the model analyzed here (except for the +1+1+ 1 defect at the origin in the inverted problem of Sec. IV), but defects may arise in other geometries Chandler and Spagnolie (2024a); Neville et al. (2024) and are likely to have an impact on the stability and dynamics. The interaction of multiple droplets suspended in an active nematic is also an open topic of interest.
@@ Acknowledgements.
D.S. acknowledges financial support from National Science Foundation Grant No. DMS-2153520.
@@
$\frac{\text{d}^{4}G_{m}}{\text{d}r^{4}}+\frac{2}{r}\frac{\text{d}^{3}G_{m}}{\text{d}r^{3}}-\frac{2m^{2}+1}{r^{2}}\frac{\text{d}^{2}G_{m}}{\text{d}r^{2}}+\frac{2m^{2}+1}{r^{3}}\frac{\text{d}G_{m}}{\text{d}r}+\frac{m^{4}-4m^{2}}{r^{4}}G_{m}=0,$ (A-1)